OpenStudy (anonymous):

write in logarithmic form: (3/4)^-1=4/3

OpenStudy (ranga):

Take log on both sides: Make use of the following logarithmic identities: log(A)^m = m * log(A) log(A/B) = log(A) - log(B)

OpenStudy (isaiah.feynman):